A sample of 100 savings accounts at a certain bank showed that 20 were overdrawn. Another sample of 125 current accounts of the bank showed that 30 were overdrawn. Test, at a 0.02 level of significance, the null hypothesis that there is no difference in the proportion of overdrawn accounts.

Answer by stanbon(75887)

A sample of 100 savings accounts at a certain bank showed that 20 were overdrawn. Another sample of 125 current accounts of the bank showed that 30 were overdrawn. Test, at a 0.02 level of significance, the null hypothesis that there is no difference in the proportion of overdrawn accounts.
-----------------------------------------
Ho: p1-p2 = 0
Ha: p1-p2 is not equal to 0
----------------------------------
p1-hat = 20/100
p2-hat = 30/125
---------------------
alpha = 2%
=========================
I ran a 2-Prop Ztest and got the following:
test stat: z = -0.7171
p-value: 0.4733
----
Since the p-value is greater than 2%, fail to reject Ho.
==============================
Cheers,
Stan H.